# Reflexive homology and involutive Hochschild homology as equivariant Loday constructions

### v3

@@ 1. Introduction
In the equivariant case we fix a finite group $G$. The starting point for a Loday construction is a $G$-commutative monoid in the sense of Hill and Hopkins [hill-hopkins]. In the setting of $G$-equivariant stable homotopy theory these are genuine $G$-commutative ring spectra whereas in the algebraic setting of Mackey functors $G$-commutative monoids are $G$-Tambara functors. Some equivariant homology theories such as the twisted cyclic nerve of Blumberg-Gerhardt-Hill-Lawson [bghl] and Hesselholt-Madsen’s Real topological Hochschild homology, $\mathsf{THR}$, [dmpr] can be identified with such equivariant Loday constructions [lrz-gloday, §7]. Here, $\mathsf{THR}$ is a homology theory for associative algebra spectra with anti-involution $A$ and we identified this in the commutative case with the Loday construction over the one-point compactification of the sign-representation, $\mathsf{THR}(A)\simeq\mathcal{L}^{C_{2}}_{S^{\sigma}}(A)$. In [lrz-gloday, Proposition 6.1], we show that for any $G$-simplicial set $X$, if we apply the functor $\underline{\pi}_{0}$ levelwise to the equivariant Loday construction of a connective genuine commutative $G$-algebra spectrum $A$ to obtain a simplicial $G$-Tambara functor,
@@
In hindsight, this identifies the Loday construction over the $C_{2}$-Burnside Tambara functor with the Loday construction relative to ¯${\underline{\mathbb{Z}}}^{c}$ under the above assumptions (see Remark 5.3). We consider the examples of $\mathbb{F}_{2}$ and $\mathbb{Z}$ with the trivial $C_{2}$-action in section 8 in order to understand what happens if we drop these assumptions. There, the homotopy groups of the Loday constructions differ both from reflexive homology and from involutive Hochschild homology.
The relationship to the Real Hochschild homology of [akgh] is more subtle: The latter takes all dihedral groups into account and for $D_{2}=C_{2}$ their definition agrees with our equivariant Loday construction. We will establish a full comparison also for the higher $D_{2m}$ with equivariant Loday constructions in future work.
@@
The proofs, however, are different: In the case of an associative ring $R$ with anti-involution the fixed point Mackey functor $\underline{R}^{\mathrm{fix}}$ is not an associative $C_{2}$-Green functor, so in particular it is not a $C_{2}$-Tambara functor. But it has the structure of a discrete $E_{\sigma}$-ring in the sense of [akgh, §6.3] and this is equivalent to a Hermitian Mackey functor in the sense of [do, Definition 1.1.1]. Real Hochschild homology [akgh, §6.4] and Real topological Hochschild homology [dmpr, Example 2.4] are defined for such objects, see also [horev, Proposition 7.1.1] for the analogous result for factorization homology of the flip circle $S^{\sigma}$, so it is not surprising that one can extend the Loday construction for $S^{\sigma}$ to fixed point Mackey functors of rings with anti-involution. We endow the $C_{2}$-Mackey norm functor of $i_{e}^{*}\underline{R}^{\mathrm{fix}}$ with the structure of an associative Green functor so that $\underline{R}^{\mathrm{fix}}$ is a bimodule over it.
@@ Acknowledgements
The first author was supported by NSF grant DMS-2004300 while writing this paper. The second author thanks the Department of Mathematics of the Indiana University Bloomington for its hospitality and support. An anonymous referee helped us to improve the paper with many insightful comments.
@@ 2. Equivariant Loday constructions
Here, $i_{H}^{*}$ restricts a $G$-Tambara functor to $H$, so for a finite $H$-set $Y$, $i_{H}^{*}{\underline{T}}(Y):={\underline{T}}(G\times_{H}Y)$. The restriction functor has the norm functor $N_{H}^{G}$ as a left adjoint. A disjoint union of $G$-sets $X,X^{\prime}$, $X\sqcup X^{\prime}$ is sent to
@@ 3. Basic results about fixed point Tambara functors
Note that for an arbitrary finite group $G$, the functor that sends a commutative $G$-ring $T$ to its fixed point $G$-Tambara functor is right adjoint to the functor that takes a $G$-Tambara functor ${\underline{R}}$ to its underlying commutative $G$-ring ${\underline{R}}(G/e)$ (see for instance [ssw, Lemma 2.9]). For $G=C_{2}$ a description of the counit map $\varrho_{T}\colon{\underline{T}}^{\mathrm{fix}}(C_{2}/e)\rightarrow T$ and the unit map $\eta_{\underline{R}}\colon{\underline{R}}\rightarrow{\underline{{\underline{R}}(C_{2}/e)}}^{\mathrm{fix}}$ of this adjunction are very straightforward: The counit $\varrho_{T}\colon{\underline{T}}^{\mathrm{fix}}(C_{2}/e)=T\rightarrow T$ is the identity map.
@@
[…]
@@
where $C_{2}$ acts on $R\otimes R$ via $\tau(a\otimes b)=\bar{b}\otimes\bar{a}$, $[a\otimes b]$ denotes the equivalence class of $a\otimes b$ in $(R\otimes R)/C_{2}$, and Tambara Reciprocity, TR, identifies $\{a+b\}\sim\{a\}+\{b\}+[a\otimes\bar{b}]$. Here $\mathsf{norm}(a\otimes b)=\{a\bar{b}\}$ and $\mathsf{tr}(a\otimes b)=[a\otimes b]$ for all $a\otimes b\in R\otimes R$, $\mathsf{res}(\{a\})=a\otimes\bar{a}$, and $\mathsf{res}([a\otimes b])=a\otimes b+\bar{b}\otimes\bar{a}$ (see [hm] for properties of the norm functor, especially Fact 4.4 in loc. cit.).
@@
and map it to ${\underline{(M\otimes N)}}^{\mathrm{fix}}$ by the identity on the free orbit and $\mathsf{res}$ on the trivial orbit, which will land in $(M\otimes N)^{C_{2}}$.
On the trivial orbit the map $\mathsf{res}$ includes $M^{C_{2}}\otimes N^{C_{2}}$ into $(M\otimes N)^{C_{2}}$ and maps $(M\otimes N)/C_{2}$ to $(M\otimes N)^{C_{2}}$ by the map $[m\otimes n]\mapsto m\otimes n+\bar{m}\otimes\bar{n}$ (which is an equivalence since $2$ is invertible in $M\otimes N$).
Then $\mathsf{res}$ maps $\underline{M}^{\mathrm{fix}}\Box\underline{N}^{\mathrm{fix}}(C_{2}/C_{2})$ surjectively onto $(M\otimes N)^{C_{2}}$ because its restriction to the second summand does, and it is injective because its restriction to the second summand is injective and Frobenius Reciprocity allows us to identify the first summand into the second one: if $2$ is invertible in $M$, any $m\in M^{C_{2}}$ is equal to $\mathsf{tr}(m/2)$ so $m\otimes n$ is identified with $[m/2\otimes n]$, and if $2$ is invertible in $N$ similarly $m\otimes n$ is identified with $[m\otimes n/2]$. As we use the identity map at the trivial level, we also get the compatibility with the Weyl group action because on both sides $C_{2}$ acts diagonally.
This is a map of Mackey functors, and, in the case where $M$ and $N$ are commutative rings with involution, also a map of Tambara functors because of the relations that $\mathsf{res}$ satisfies with respect to the other structure maps. ∎
@@
The formula for the box product (see e.g. [hm, Definition 3.1]) gives
@@
and again we propose to send it to ${\underline{(R\otimes R\otimes M)}}^{\mathrm{fix}}$ by taking the identity on the free orbit and $\mathsf{res}$ on the trivial orbit, which will land in $(R\otimes R\otimes M)^{C_{2}}$.
Again, since $2$ is invertible the restriction of $\mathsf{res}$ to $(R\otimes R\otimes M)/C_{2}$, which sends $[a\otimes b\otimes m]\mapsto a\otimes b\otimes m+\bar{b}\otimes\bar{a}\otimes\bar{m}$ is an isomorphism $(R\otimes R\otimes M)/C_{2}\to(R\otimes R\otimes M)^{C_{2}}$. Frobenius Reciprocity identifies $\mathsf{tr}(a\otimes b)\otimes m$ with $[a\otimes b\otimes\mathsf{res}(m)]\in(R\otimes R\otimes M)/C_{2}$ for all $a,b\in R$, $m\in M$. It also identifies $\{a\}\otimes\mathsf{tr}(m)$ with $[a\otimes\bar{a}\otimes m]\in(R\otimes R\otimes M)/C_{2}$ for all $a\in R$, $m\in M$, and since $2$ is invertible, $\mathsf{tr}(M)=M^{C_{2}}$ (any $m\in M^{C_{2}}$ is equal to $\mathsf{tr}(m/2)$). So in fact,
@@
and our map between $(N^{C_{2}}_{e}R)\Box\underline{M}^{\mathrm{fix}}$ and ${\underline{(R\otimes R\otimes M)}}^{\mathrm{fix}}$ is an isomorphism at both levels. It preserves the Mackey and Tambara structure maps because of the relations $\mathsf{res}$ satisfies with respect to the structure maps. ∎
@@ 4. Working relative to a commutative ground ring
The purpose of this section is to relate the relative Loday construction of $C_{2}$-fixed point Tambara functors to the fixed point Tambara functor of the non-equivariant relative Loday construction. To that end we prove two crucial auxiliarly results.
@@
[…]
@@
[…]
@@ 5. Identifying $\mathcal{L}_{S^{\sigma}}^{C_{2}}(\underline{R}^{\mathrm{fix}})$
In this section we will continue to work with the cyclic group of order $2$, $C_{2}=\langle\tau\mid\tau^{2}=e\rangle$, and we will consider the $C_{2}$-simplicial set $S^{\sigma}$ which is the one-point compactification of the real sign-representation,
@@
where the $C_{2}$-action flips the two arcs.
@@
Note that in contrast to Proposition 4.1 $N^{C_{2}}_{e}i^{*}_{e}\underline{R}^{\mathrm{fix}}$ is not isomorphic to ${\underline{(R\otimes R)}}^{\mathrm{fix}}$, even in very simple cases! For example, for $R=\mathbb{Z}$ with the trivial $C_{2}$-action, ${\underline{(R\otimes R)}}^{\mathrm{fix}}$ is just ¯$\underline{\mathbb{Z}}^{c}$ with respect to the constant action, while $N^{C_{2}}_{e}i^{*}_{e}\underline{\mathbb{Z}}^{c}$ is the $C_{2}$-Burnside Tambara functor (see for instance [lrz-gloday, (5.1)]). We need an outer copy of ${\underline{R}}^{\mathrm{fix}}$ in Theorem 5.1 as a catalyst in order to achieve the desired simplification.
@@
It is important to remember that the equivariant Loday construction $\mathcal{L}_{X}^{C_{2}}(\underline{R}^{\mathrm{fix}})$ is not the Loday construction relative to ¯${\underline{\mathbb{Z}}}^{c}$, but rather the Loday construction relative to the $C_{2}$-Burnside Tambara functor, and these are different. For example, taking the relative norm-restriction term from (4.1) $N_{e}^{C_{2},{\underline{\mathbb{Z}}}^{c}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c}$ gives ¯${\underline{\mathbb{Z}}}^{c}$, whereas taking $N_{e}^{C_{2}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c}$ gives the $C_{2}$-Burnside Tambara functor as explained for instance in [lrz-gloday, (5.1)].
@@
where $C_{2}$ acts on $R\otimes R$ as in (5.2). Note that in the bar construction $B(R,R\otimes R,R)$ the ground ring is the ring of integers whereas for the Loday construction we work relative to the Burnside Tambara functor. Similarly, Theorem 4.3 implies that for $k=\mathbb{Z}$ we also obtain that $\mathcal{L}^{C_{2},{\underline{\mathbb{Z}}}^{c}}_{S^{\sigma}}({\underline{R}}^{\mathrm{fix}})$ can be identified with ${\underline{B^{\mathbb{Z}}(R,R\otimes R,R)}}^{\mathrm{fix}}$ and therefore in hindsight we obtain that in this case the Loday construction relative to the Burnside Tambara functor agrees with the one relative to ¯${\underline{\mathbb{Z}}}^{c}$.

### v4

@@ 1. Introduction
In the equivariant case we fix a finite group $G$. The starting point for a Loday construction is a $G$-commutative monoid in the sense of Hill and Hopkins [hill-hopkins]. In the setting of $G$-equivariant stable homotopy theory these are genuine $G$-commutative ring spectra whereas in the algebraic setting of Mackey functors $G$-commutative monoids are $G$-Tambara functors. Some equivariant homology theories such as the twisted cyclic nerve of Blumberg-Gerhardt-Hill-Lawson [bghl] and Hesselholt-Madsen’s Real topological Hochschild homology, $\mathsf{THR}$, [dmpr] can be identified with such equivariant Loday constructions [lrz-gloday, §7]. Here, $\mathsf{THR}$ is a homology theory for associative algebra spectra with anti-involution $A$ and we identified this in the commutative case with the Loday construction over the one-point compactification of the sign-representation, $\mathsf{THR}(A)\simeq\mathcal{L}^{C_{2}}_{S^{\sigma}}(A)$. In the following we will often refer to $S^{\sigma}$ as the flip circle. In [lrz-gloday, Proposition 6.1], we show that for any $G$-simplicial set $X$, if we apply the functor $\underline{\pi}_{0}$ levelwise to the equivariant Loday construction of a connective genuine commutative $G$-algebra spectrum $A$ to obtain a simplicial $G$-Tambara functor,
@@
In hindsight, this identifies the Loday construction over the $C_{2}$-Burnside Tambara functor with the Loday construction relative to ${\underline{\mathbb{Z}}}^{c}$ under the above assumptions (see Remark 5.3). We consider the examples of $\mathbb{F}_{2}$ and $\mathbb{Z}$ with the trivial $C_{2}$-action in section 8 in order to understand what happens if we drop these assumptions. There, the homotopy groups of the Loday constructions differ both from reflexive homology and from involutive Hochschild homology.
The relationship to the Real Hochschild homology of [akgh] is more subtle: The latter takes all dihedral groups into account and for $D_{2}=C_{2}$ their definition agrees with our equivariant Loday construction. We will establish a full comparison also for the higher $D_{2m}$ with equivariant Loday constructions in future work with Foling Zou (in preparation).
@@
The proofs, however, are different: In the case of an associative ring $R$ with anti-involution the fixed point Mackey functor $\underline{R}^{\mathrm{fix}}$ is not an associative $C_{2}$-Green functor, so in particular it is not a $C_{2}$-Tambara functor. But it has the structure of a discrete $E_{\sigma}$-ring in the sense of [akgh, §6.3] and it is also a Hermitian Mackey functor in the sense of [do, Definition 1.1.1]. Real Hochschild homology [akgh, §6.4] and Real topological Hochschild homology [dmpr, Example 2.4] are defined for such objects, see also [horev, Proposition 7.1.1] for the analogous result for factorization homology of the flip circle $S^{\sigma}$, so it is not surprising that one can extend the Loday construction for $S^{\sigma}$ to fixed point Mackey functors of rings with anti-involution. We endow the $C_{2}$-Mackey norm functor of $i_{e}^{*}\underline{R}^{\mathrm{fix}}$ with the structure of an associative Green functor so that $\underline{R}^{\mathrm{fix}}$ is a bimodule over it.
@@ Acknowledgements
The first author was supported by NSF grant DMS-2004300 at the beginning of working on this paper, and by a grant from the Simons Foundation (#359565, Ayelet Lindenstrauss) afterwards. The second author thanks the Department of Mathematics of the Indiana University Bloomington for its hospitality and support. An anonymous referee helped us to improve the paper with many insightful comments.
@@ 2. Equivariant Loday constructions
Here, $i_{H}^{*}$ restricts a $G$-Tambara functor to $H$, so for a finite $H$-set $Y$, $i_{H}^{*}{\underline{T}}(Y):={\underline{T}}(G\times_{H}Y)$. The restriction functor has the norm functor $N_{H}^{G}$ as a left adjoint. A disjoint union of $G$-sets $^{\prime}, XXX\sqcup X^{\prime}$ is sent to
@@ 3. Basic results about fixed point Tambara functors
Remark 3.1.
Note that for an arbitrary finite group $G$, the functor that sends a commutative $G$-ring $T$ to its fixed point $G$-Tambara functor is right adjoint to the functor that takes a $G$-Tambara functor $\underline{R}$} to its underlying commutative $G$-ring ${\underline{R}}(G/e)$ (see for instance [ssw, Lemma 2.9]). This is completely analogous to the situation in Mackey functors, where the functor that sends an abelian group $A$ with $G$-action to its fixed point $G$-Mackey functor ${\underline{A}}^{\mathrm{fix}}$ is right adjoint to the functor that takes a $G$-Mackey functor $\underline{M}$} and sends it to the abelian group with $G$-action ${\underline{M}}(G/e)$.
For $G=C_{2}$ a description of the counit map $\varrho_{T}\colon{\underline{T}}^{\mathrm{fix}}(C_{2}/e)\rightarrow T$ and the unit map $\eta_{\underline{R}}\colon{\underline{R}}\rightarrow{\underline{{\underline{R}}(C_{2}/e)}}^{\mathrm{fix}}$ of this adjunction are very straightforward (both in the Mackey case and in the Tambara case): The counit $\varrho_{T}\colon{\underline{T}}^{\mathrm{fix}}(C_{2}/e)=T\rightarrow T$ is the identity map.
@@
${{\underline{k}}^{c}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\underline{i_{T}}}}$$\scriptstyle{{\underline{i_{R}}}}$${\underline{R}^{\mathrm{fix}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{{\underline{f}}}$${{\underline{T}}^{\mathrm{fix}}}$
@@
where $C_{2}$ acts on $R\otimes R$ via $\tau(a\otimes b)=\bar{b}\otimes\bar{a}$, $a\otimes b]$ denotes the equivalence class of $a\otimes b$ in $(R\otimes R)/C_{2}$, and Tambara Reciprocity, TR, identifies $\{a+b\}\sim\{a\}+\{b\}+[a\otimes\bar{b}]$. Here $\mathsf{norm}(a\otimes b)=\{a\bar{b}\}$ and $\mathsf{tr}(a\otimes b)=[a\otimes b]$ for all $a\otimes b\in R\otimes R$, $\mathsf{res}(\{a\})=a\otimes\bar{a}$, and $\mathsf{res}([a\otimes b])=a\otimes b+\bar{b}\otimes\bar{a}$ (see [hm] for properties of the norm functor, especially Fact 4.4 in loc. cit.).
@@
and we map it to ${\underline{(M\otimes N)}}^{\mathrm{fix}}$ by using the Mackey or Tambara functor map corresponding to the identity map on the free orbit $C_{2}/e$ via the adjunction in Remark 3.1. Then our map will consist of the identity on the free orbit and $\mathsf{res}$ on the trivial orbit, which will land in $(M\otimes N)^{C_{2}}$.
On the trivial orbit the map $\mathsf{res}$ includes $M^{C_{2}}\otimes N^{C_{2}}$ into $(M\otimes N)^{C_{2}}$ and maps $(M\otimes N)/C_{2}$ to $(M\otimes N)^{C_{2}}$ by the map $m\otimes n]\mapsto m\otimes n+\bar{m}\otimes\bar{n}$ (which is an equivalence since $2$ is invertible in $M\otimes N$).
Then $\mathsf{res}$ maps $\underline{M}^{\mathrm{fix}}\Box\underline{N}^{\mathrm{fix}}(C_{2}/C_{2})$ surjectively onto $(M\otimes N)^{C_{2}}$ because its restriction to the second summand does, and it is injective because its restriction to the second summand is injective and Frobenius Reciprocity allows us to identify the first summand into the second one: if $2$ is invertible in $M$, any $m\in M^{C_{2}}$ is equal to $\mathsf{tr}(m/2)$ so $m\otimes n$ is identified with $m/2\otimes n]$, and if $2$ is invertible in $N$ similarly $m\otimes n$ is identified with $m\otimes n/2]$. ∎
@@
Applying the formula for the box product ([hm, Definition 3.1]) to the formula for the norm yields
@@
and again we send it to ${\underline{(R\otimes R\otimes M)}}^{\mathrm{fix}}$ by the Mackey or Tambara functor map that corresponds via the adjunction of Remark 3.1 to the identity on the free orbit. So the resulting map is the identity on the free orbit and $\mathsf{res}$ on the trivial orbit, which will land in $(R\otimes R\otimes M)^{C_{2}}$.
Again, since $2$ is invertible the restriction of $\mathsf{res}$ to $(R\otimes R\otimes M)/C_{2}$, which sends $a\otimes b\otimes m]\mapsto a\otimes b\otimes m+\bar{b}\otimes\bar{a}\otimes\bar{m}$ is an isomorphism $(R\otimes R\otimes M)/C_{2}\to(R\otimes R\otimes M)^{C_{2}}$. Frobenius Reciprocity identifies $\mathsf{tr}(a\otimes b)\otimes m$ with $a\otimes b\otimes\mathsf{res}(m)]\in(R\otimes R\otimes M)/C_{2}$ for all $a,b\in R$, $m\in M^{C_{2}}$. It also identifies $\{a\}\otimes\mathsf{tr}(m)$ with $a\otimes\bar{a}\otimes m]\in(R\otimes R\otimes M)/C_{2}$ for all $a\in R$, $m\in M$, and since $2$ is invertible, $\mathsf{tr}(M)=M^{C_{2}}$ (any $m\in M^{C_{2}}$ is equal to $\mathsf{tr}(m/2)$). So in fact,
@@
and our map between $(N^{C_{2}}_{e}R)\Box\underline{M}^{\mathrm{fix}}$ and ${\underline{(R\otimes R\otimes M)}}^{\mathrm{fix}}$ is an isomorphism at both levels. ∎
@@ 4. Working relative to a commutative ground ring
The purpose of this section is to relate the relative Loday construction of $C_{2}$-fixed point Tambara functors to the fixed point Tambara functor of the non-equivariant relative Loday construction. To that end we prove two crucial auxiliary results.
@@
${N_{e}^{C_{2}}i_{e}^{*}({\underline{R}}^{\mathrm{fix}})\Box N_{e}^{C_{2}}i_{e}^{*}{\underline{k}}^{c}\Box{\underline{k}}^{c}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\nu\Box\mathrm{id}}$$\scriptstyle{\mathrm{id}\Box\nu^{\prime}}$${N_{e}^{C_{2}}i_{e}^{*}({\underline{R}}^{\mathrm{fix}})\Box{\underline{k}}^{c}}$
@@
where $cC_{2}\text{-rings}$ denotes the category of commutative $C_{2}$-rings. Therefore, if we restrict to the above full subcategory, taking the fixed point Tambara functor is left adjoint, hence preserves coequalizers.
@@ 5. Identifying $\mathcal{L}_{S^{\sigma}}^{C_{2}}(\underline{R}^{\mathrm{fix}})$
In this section we will continue to work with the cyclic group of order $2$, $C_{2}=\langle\tau\mid\tau^{2}=e\rangle$, and we will consider the $C_{2}$-simplicial set $S^{\sigma}$ which models the one-point compactification of the real sign-representation,
@@
where the $C_{2}$-action flips the two arcs. We will call $S^{\sigma}$ with this action the flip circle.
@@
Note that in contrast to Proposition 4.1 $N^{C_{2}}_{e}i^{*}_{e}\underline{R}^{\mathrm{fix}}$ is not isomorphic to ${\underline{(R\otimes R)}}^{\mathrm{fix}}$, even in very simple cases! For example, for $R=\mathbb{Z}$ with the trivial $C_{2}$-action, ${\underline{(R\otimes R)}}^{\mathrm{fix}}$ is just $\underline{\mathbb{Z}}^{c}$ with respect to the constant action, while $N^{C_{2}}_{e}i^{*}_{e}\underline{\mathbb{Z}}^{c}$ is the $C_{2}$-Burnside Tambara functor (see for instance [lrz-gloday, (5.1)]). We need an outer copy of ${\underline{R}}^{\mathrm{fix}}$ in Theorem 5.1 as a catalyst in order to achieve the desired simplification.
@@
It is important to remember that the equivariant Loday construction $\mathcal{L}_{X}^{C_{2}}(\underline{R}^{\mathrm{fix}})$ is not the Loday construction relative to ${\underline{\mathbb{Z}}}^{c}$, but rather the Loday construction relative to the $C_{2}$-Burnside Tambara functor, and these are different. For example, taking the relative norm-restriction term from (4.1) $N_{e}^{C_{2},{\underline{\mathbb{Z}}}^{c}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c}$ gives ${\underline{\mathbb{Z}}}^{c}$, whereas taking $N_{e}^{C_{2}}i_{e}^{*}{\underline{\mathbb{Z}}}^{c}$ gives the $C_{2}$-Burnside Tambara functor.
@@
where $C_{2}$ acts on $R\otimes R$ as in (5.2). Note that in the bar construction $B(R,R\otimes R,R)$ the ground ring is the ring of integers whereas for the Loday construction we work relative to the Burnside Tambara functor. Similarly, Theorem 4.3 implies that for $k=\mathbb{Z}$ we also obtain that $\mathcal{L}^{C_{2},{\underline{\mathbb{Z}}}^{c}}_{S^{\sigma}}({\underline{R}}^{\mathrm{fix}})$ can be identified with ${\underline{B^{\mathbb{Z}}(R,R\otimes R,R)}}^{\mathrm{fix}}$ and therefore in hindsight we obtain that in this case the Loday construction relative to the Burnside Tambara functor agrees with the one relative to ${\underline{\mathbb{Z}}}^{c}$.